OpenStudy (anonymous):

You have a solution made of 300 grams of ethanol (C2H5OH) and 500 grams of water. What is the mole fraction of ethanol in this solution? (Points : 1) 0.19 M 0.81 M 0.19 0.81

OpenStudy (anonymous):

If you want to make 0.5 L of a 0.01 M solution of bromine (Br2) in water, how much bromine would you need? (Points : 1) 0.8 grams 0.8 moles 1.6 grams 1.6 moles Question 2.2. Which of the following statements is the correct definition of molarity? (Points : 1) number of moles of solvent/liter solution number of grams of solvent/liter solution number of moles of solute/liter solution number of grams of solute/liter solution

OpenStudy (anonymous):

the last one is \[M=\frac{ moles of solute }{ Liters of solution }\]

OpenStudy (anonymous):

so its C

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

for the other two, its just about the same concepts. do you know how to use the molar mass to convert to moles?

OpenStudy (anonymous):

no i dont

OpenStudy (anonymous):

okay well this should help: \[grams of substance*\frac{ moles of substance }{ molar mass of substance }\]

OpenStudy (anonymous):

do you know what the molar mass is?

OpenStudy (anonymous):

the mass of a substancd

OpenStudy (anonymous):

the mass of a substance per one mole. you get the molar mass by using the periodic table of elements

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

if you use the equation i gave you to convert, then those other two problems should be a breeze.

OpenStudy (anonymous):

alright i will if i need any help ill ask :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I dont know the first one but for the second one i got B is that righ?

OpenStudy (anonymous):

very close but not quite. \[M=\frac{ mol solute }{ L solution }\]

OpenStudy (anonymous):

so....\[0.01=\frac{ x moles }{ .5 L }\]

OpenStudy (anonymous):

x moles = 0.005 moles Br2

OpenStudy (anonymous):

but that isnt one of the options so we need to go farther. (0.005 mol)(159.808g/1 mol) "mol" cancels and we are left with 0.8 grams Br2

OpenStudy (anonymous):

Ohhh okay. what about the other one

OpenStudy (anonymous):

use the molar mass to convert both substances to moles first.

OpenStudy (anonymous):

Ethanol: (500g)(1 mol/46.07g) 10.85 mol ethanol

OpenStudy (anonymous):

can you do the other one?

OpenStudy (anonymous):

no i dont understand it

OpenStudy (anonymous):

its thesame thing that i just did. with ethanol

OpenStudy (anonymous):

with the what? sorry im confused

OpenStudy (anonymous):

sometimes its easier to see it written.... does this help?

OpenStudy (anonymous):

yes i skipped over the water part thats why i didnt understand

OpenStudy (anonymous):

okay. so do you know what a mole fraction is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\frac{ moles-of-substance }{ total-number-of-moles }\]

OpenStudy (anonymous):

so... we want this: \[\frac{ moles-of-ethanol }{ (moles-of-ethanol)+(moles-of-water) }\]

OpenStudy (anonymous):

so the answer is .19m

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Yayyy thank you

OpenStudy (anonymous):

do you understand it? like maybe with a bit of practice, you could do it on your own next time?

OpenStudy (anonymous):

yea i think so

OpenStudy (anonymous):

good. :)